### v3

@@ 5 Optimal Design
Recall that $d$ is the dimension of the spatial region, $\xi_{n}$ is the density of the region (number of units per unit volume/area), and $\gamma$ is the speed at which interference decays with distance. By Theorem 2,
@@
optimizes the rate of convergence of $\hat{\theta}$. This formalizes how domain knowledge of interference $\gamma$ informs the design of clusters. The right-hand side is increasing in $\gamma$ since less interference means we can achieve the same level of bias reduction with more clusters. It is decreasing in density $\xi_{n}$ because having more units in a given area effectively corresponds to greater interference (an infectious disease may spread more easily).
Choosing the number of clusters according to (8) corresponds to balancing the asymptotic orders of the bias and variance of $\hat{\theta}$. However, as discussed in § 4, validity of the CI (5) requires the bias to be of smaller order than the variance. This requires choosing $k$ smaller than the optimal choice (8), which is what we called an undersmoothed design in § 4.
We next provide a recommendation for choosing an undersmoothed $k$. Let $\mathcal{V}$ denote the area or volume of the study region containing $\mathcal{N}_{n}$. As discussed below, $\mathcal{V}$ is of order $n/\xi_{n}$, so the rate-optimal formula (8) can be rewritten as $k\sim\mathcal{V}^{\frac{2\gamma}{2\gamma+d}}$. While one could choose $k$ equal to the right-hand side, the problem is that this choice is extremely sensitive to the unit of length used to measure geographic distance. Switching from square kilometers to square millimeters would dramatically increase $\mathcal{V}$.
Our observation is that both the unit of length and $\gamma$ determine the speed at which interference decays with distance. For any given choice of $\gamma$, the rate of decay is significantly faster if distance is measured in millimeters compared to kilometers, for example. One should therefore tailor both $\gamma$ and the unit of length to domain knowledge of interference and select $k$ as follows.
@@
The next subsection provides examples of calibrating $k$ under the conservative choice $\tilde{\gamma}=d$. The last subsection discusses how to obtain less conservative estimates.
@@ 5.1 Empirical Examples
Sur et al. (2009) study an area of roughly 1.2 by 0.7 km containing 38k units. To compute our suggested number of clusters (9), we need to select $\tilde{\gamma}$ and the unit of length. Suppose we choose the conservative bound $\tilde{\gamma}=d=2$, so that interference is assumed to decay like $r^{-2}$ or faster with each unit of length $r$. If we take 35m to be the unit of length, then for units $i,j,k$ such that $\rho(i,j)=35$m and $\rho(i,k)=70$m, the extent to which $k$’s treatment affects $i$ is less than $1/4$th ($0.252^{-\tilde{\gamma}}=$ 0.25) as small as the extent to which $j$’s affects $i$. This is with only a 35m difference in distance. For this unit of length, (9) yields $k=78$. Hence, these are the assumptions on interference that justify the authors’ choice of $k=80$, originally determined by a conventional power analysis based on a partial interference model.
@@
Next consider Homan et al. (2016) whose trial region is substantially larger at roughly 12 by 4 km with nearly the same number of units (34k). Due to the lower density, there is less bias from interference, so a choice of $k=81$ can be justified under weaker assumptions on interference. Specifically, if $\tilde{\gamma}=2$ but the unit of length is now in 250m increments, then (9) results in $k=84$. For context, the Anopheles mosquito that is the subject of their trial typically does not fly more than 2 km from their breeding grounds (CDC, 2023).
@@ 5.2 Bounding Interference
Any optimal design generally requires prior information on certain population parameters. The standard power analysis for CRTs assumes partial interference and requires knowledge of the intracluster correlation coefficient, a measure of within-cluster dependence in potential outcomes (e.g. Baird et al., 2018; Hemming et al., 2017). Our results suggest that, if cross-cluster interference is of first-order importance, the focus of attention should instead be determining $\gamma$. As illustrated in the previous subsection, the conservative choice $\gamma=d$ can result in relatively small values of $k$ because it allows for a greater degree of interference, so to the extent that one can justify stronger assumptions on interference, that is, values of $\gamma$ that are larger than $d$ for a given unit of length, this would improve asymptotic power.
In the context of infectious disease trials, Halloran et al. (2017) argue that CRT design should be informed by simulating models of disease transmission. Several papers utilize parametric models and simulation methods to estimate or bound contamination bias. Alexander et al. (2020) and Jarvis et al. (2019) use spatial models to provide evidence of contamination in prior CRTs. Multerer et al. (2021) employ models of disease transmission for a similar purpose. Our theory provides a precise way in which modeling can inform design, namely by providing estimates of $\gamma$. This relates to Staples et al. (2015) who show how to estimate a different measure of cross-cluster interference to assess the degree to which the conventional power analysis overstates trial power.
@@ 6 Simulation Study
Let $\{\tilde{\varepsilon}_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(-0.5,1)$, $\{\beta_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(2,1)$, and $\{\gamma_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(1,1)$ be independent and drawn independently of locations. For $G_{ij}=\bm{1}\{\rho(i,j)\leq 1\}/\sum_{k\in\mathcal{N}_{n}}\bm{1}\{\rho(i,k)\leq 1\}$, let $\varepsilon_{i}=\tilde{\varepsilon}_{i}+\sum_{j\in\mathcal{N}_{n}}G_{ij}\tilde{\varepsilon}_{j}$, which serves to generate spatial autocorrelation in errors. For $w_{ij}=\min\{\rho(i,j)^{-5},1\}$, we generate outcomes according to
$Y_{i}=\sum_{j\in\mathcal{N}_{n}}w_{ij}D_{j}\beta_{j}+\sum_{j\in\mathcal{N}_{n}}w_{ij}D_{i}D_{j}\gamma_{j}+\varepsilon_{i}$
@@
We present results for the indirect and total effects using 5000 simulation draws, where within each simulation, we redraw potential outcomes and recompute the design-based estimand. In Table 1, the “Spatial Interference” columns correspond to the outcome model described above, whereas the “Partial Interference” columns redefine $w_{ij}=0$ if $i,j$ lie in different clusters to eliminate cross-cluster interference. The “CI” rows report the coverage of 95-percent CIs using the indicated standard errors. For our estimator, the “SE” row corresponds to standard errors obtained from our variance estimator (4), while for difference in means, it corresponds to conventional cluster-robust SEs. The “SE” rows are the true superpopulation standard errors obtained by taking the standard deviation of the estimator across the simulation draws. Hence SE should be consistent for SE rather than conservative, whereas the CIs are asymptotically conservative in terms of coverage of the design-based estimand. Finally, row “Excl” is the percentage of units that are not well surrounded.
@@
Table 2 presents different values of $r_{n}$ under spatial interference to assess the robustness of our recommended choice (2) and illustrate a bias-variance trade-off in the choice of $r_{n}$. We replace $r_{n}$ with $c$ times (2) for $c\in\{0.8,1.2\}$. The $c=0.8$ columns show that this results in about half the proportion of excluded units relative to Table 1. As a result, the bias is higher, resulting in undercoverage. The $c=1.2$ columns show that the proportion of excluded units a little less than doubles. The bias is lower and as a result, the probability of coverage is higher. On the other hand, the variance increases, as can be seen in the standard error columns.
@@ 7 Empirical Application
We use our proposed estimator to reanalyze data from the unconditional cash transfer experiment conducted by Egger et al. (2022) mentioned in Example 2. In their experiment, households eligible for the transfer (which is the treatment) live in homes with thatched roofs, which is a proxy for poverty. Houses are grouped in villages, which are grouped in “sublocations.” The CRT randomizes sublocations (the clusters) into treatment with probability $q=0.5$. Within treatment (control) sublocations, villages are randomized into treatment with probability $p_{1}=2/3$ ($p_{0}=1/3$). Within treated villages, all eligible households receive cash transfers totaling 1000 USD, which is about 75 percent of average annual household spending. Sublocations contain on average 7.8 villages with a standard deviation of 3.9.
@@
However, the authors note the potential for interference across sublocations (their quote in Example 2). For this reason, their preferred specification is the following instrumental variables (IV) regression. The main regressors are the amount of cash transferred per capita to the household’s village and the amount transferred to neighboring villages within a band of $r-2$ to $r$ km from the ego’s village for a range of $r$ values. The corresponding instruments are respectively an indicator for the ego’s village being treated and the share of eligible households in the band assigned to treatment. They use a BIC criterion to select the maximum range of $r$, which is 2 km. Using these estimates, the authors compute a model-assisted estimate of the total effect $\theta^{*}(1,0,2/3,0)$. This compares saturation levels of two-thirds and zero, which is nonparametrically unidentified since control villages have a saturation level of one-third. For this reason, we report our estimates of the total effect $\theta^{*}(1,0,2/3,1/3)$.
Table 3 reports the results for a subset of the outcomes and Table SA.4.1 in the supplementary appendix reports the remainder. Both tables choose $r_{n}$ in $\hat{\theta}$ according to (2), which results in $r_{n}=1.6$ and 39.66 percent of units not well surrounded. In § SA.4 we describe how we construct cluster radii $R_{j}$ used in this formula. The $\hat{\theta}^{+}$ columns correspond to difference-in-means estimates. The design-based estimates of the direct and total effects are quite similar in magnitude to the RF and IV estimates, respectively, despite the distinctions outlined above. Decomposing the total effect into direct and indirect effects, we find that the latter are substantially smaller in magnitude with large standard errors. Compared to difference in means, our estimates tend to be larger in magnitude, which is the impact on bias, but so are the standard errors, which is the impact on variance.
@@
To estimate spillover effects, Egger et al. (2022) rerun their IV specification using only non-eligible households, which did not receive any transfers. In Table SA.4.3 of the supplementary appendix, we compare this with design-based estimates of the overall effect $\theta^{*}(\varnothing,\varnothing,2/3,1/3)$ on the population of non-eligibles. The effect sizes of our estimators and those of difference in means are fairly similar in magnitude to their IV results, but the standard errors are large.
@@ 8 Conclusion
This paper studies inference on standard CRT estimands using data from two-stage randomized designs. We provide a tractable theoretical framework for relaxing the partial interference assumption and allowing for cross-cluster interference. To reduce bias from interference at the analysis stage, we provide in § 3.2 an estimator $\hat{\theta}$ that improves upon the fried-egg design. This can be used as a robustness check for difference in means. To reduce bias at the design stage, we propose choosing the number of clusters $k$ according to (9) instead of the standard power analysis that assumes partial interference. If the design parameters include the clusters themselves, we suggest automating cluster construction using $k$-medoids (§ 2.2).
Under the conventional superpopulation, partial interference framework, prior knowledge of the intracluster correlation coefficient (ICC) is necessary for power calculations to choose $k$. Under our design-based framework, power is instead determined by $\gamma$, the speed at which interference decays with distance. Thus, our optimal formula for $k$ requires the experimenter to take a stance on $\gamma$ rather than the ICC. Absent domain knowledge, one can select the conservative choice $\gamma=d$, but we discuss in § 5.2 how to obtain less conservative estimates from prior data such as pilot studies. The determination of $\gamma$ is an important topic for future research.
@@
Whereas $k$-means allows centroids to be in $\mathcal{X}$, $k$-medoids constrains the centroids to be in the data $\mathcal{N}_{n}\subseteq\mathcal{X}$, but in practice the algorithms tend to produce similar output. Our theoretical results pertain to the implementation in Algorithm SA.1.1, the “partitioning around medoids” algorithm. This is simple to understand, though it has complexity $O(k(n-k)^{2})$ compared to an $O(n^{2})$ runtime under the fastest known implementation (Schubert and Rousseeuw, 2021).
@@
Call the density-homogeneous subregions $\mathcal{S}_{1},\ldots,\mathcal{S}_{m}$. For each $\mathcal{S}_{j}$ the researcher can compute the optimal number of clusters $k_{j}$ given in (9), subdivide $\mathcal{S}_{j}$ into $k_{j}$ clusters $C_{j1},\ldots,C_{jk_{j}}$, say using $k$-medoids, and cluster-randomize across the collection of all clusters $\{C_{j\ell}\colon\ell=1,\ldots k_{j},j=1,\ldots,m\}$. Lastly, we recommend modifying the formula for $r_{n}$ since cluster radii may vary substantially across subregions with differing densities. In the definition of $T_{ti}$, replace $r_{n}$ with $r_{n}(i)$, defined as half the median cluster radius among clusters in the subregion containing $i$.
@@
This section relates our variance estimator to prior work in the literature. Define $\bm{A}(1)$ ($\bm{A}(2)$) as the symmetric, $n\times n$ matrix with $ij$th entry $A_{ij}(1)$ ($A_{ij}(2)$), and recall that $\hat{\sigma}^{2}(2)$ is the cluster-robust variance estimator, while $\hat{\sigma}^{2}(1)$ is analogous to the Leung (2022b) variance estimator. To establish asymptotic conservativeness of $\hat{\sigma}^{2}(1)$, Leung (2022b) requires conditions on the superpopulation, which find is unnecessary.
The proof of Theorem 5 shows that the cluster-robust variance estimator can be asymptotically decomposed as $\hat{\sigma}^{2}(2)=\sigma_{n}^{2}+\mathcal{B}_{n}+o_{p}(1)$ where
@@
The proof further shows that $\hat{\sigma}^{2}(1)=\sigma_{n}^{2}+\mathcal{B}_{n}(1)+o_{p}(1)$ where $\mathcal{B}_{n}(1)$ is obtained by replacing $\bm{A}(2)$ with $\bm{A}(1)$ in (SA.3.1). This entails additional nonzero off-diagonal elements to $\bm{A}(2)$, so $\bm{A}(1)$ is not block-diagonal and hence not guaranteed to be positive semidefinite. Accordingly, Leung imposes additional conditions to show that $\mathcal{B}_{n}(1)\stackrel{{\scriptstyle p}}{{\longrightarrow}}c\geq 0$ in the superpopulation. We find, however, that $\mathcal{B}_{n}(1)=\mathcal{B}_{n}+o_{p}(1)$ without any conditions on the superpopulation, so in fact $\hat{\sigma}^{2}(1)$ is asymptotically conservative in a purely design-based setup. Faridani and Niehaus (2023) provide a different approach to conservative inference for the GATE based on bounding $\mathcal{B}_{n}$.
The positive-semidefiniteness of the “kernel” $\bm{A}(2)$ results in both a non-negative variance estimator for any $n$ and asymptotic conservativeness. This mirrors the corresponding insight for HAC variance estimators, that positive semidefinite kernels ensure conservativeness, which was first pointed out in the literature by Leung (2019) and has since been exploited by other papers. In his case of network-dependent data, the difficulty with using positive-semidefinite HAC kernels is that they are sloped and tend to severely over-reject in finite samples, which is why Leung (2022a) recommends use of the uniform kernel despite the fact that it is not positive semidefinite. For spatial data, however, a variety of positive semidefinite HAC kernels exist, and these tend to have less severe issues with over-rejection relative to the spatial case.
Hudgens and Halloran (2008) propose variance estimators for difference-in-means type estimators under partial interference. Their theory relies on an additional stratified interference assumption, which says that potential outcomes only depend on own treatment assignment and the proportion of treated units in the ego’s cluster, in which case $Y_{i}=Y_{i}(D_{i},p)$ (since they assume complete randomization). Because $p$ is a constant, the dependence structure is the same as under no interference since outcomes within a cluster are only correlated if treatment assignments are. The variance estimators proposed by Hudgens and Halloran (2008) heavily rely on this structure. In contrast, we do not impose partial, let alone stratified, interference, so our setting features both within- and cross-cluster dependence in outcomes. It is therefore critical to account for additional covariance terms that are absent in the Hudgens and Halloran (2008) variance formula to avoid anti-conservativeness. Our estimator does so through the terms involving $A_{ij}(u)$ for $i\neq j$.
@@
We bound the discrepancy between $\tilde{Z}_{i}$ and $X_{i}^{r}$. For any $r\geq r_{n}$ and $t\in\{0,1\}$, $T_{ti}$ is measurable with respect to the $\sigma$-algebra generated by $\mathcal{F}_{i}(r)$. Then by Assumption 3, for any $q>0$, $r\geq r_{n}$, and $t\in\{0,1\}$,
@@
because $T_{ti}^{+}Y_{i}/p_{t}^{+}$ is uniformly bounded by Assumption 1 and Lemma SA.5.3, and $p_{t}^{+}-\hat{p}_{t}^{+}\precsim k_{n}^{-1/2}$ by Lemma SA.5.6. It suffices to derive the rate of convergence of $\tilde{\theta}^{+}$.
By the argument used to bound (SA.6.4) in the proof of Theorem 2, $\text{Var}(\tilde{\theta}^{+})\precsim k_{n}^{-1}$. Then it remains to bound the bias, that is, to show that
@@
$\max_{j}\lvert\mathcal{B}(C_{j})\rvert\leq\sum_{\ell=0}^{\lceil\alpha_{n}\rceil}\lvert\mathcal{N}(m_{j},L_{n}+\ell+1)\backslash\mathcal{N}(m_{j},L_{n}+\ell)\rvert\precsim\alpha_{n}\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}.$ (SA.6.15)

### v4

@@ 5 Optimal Design
Recall that $d$ is the dimension of the spatial region, $\xi_{n}$ is the density of the region (number of units per unit volume/area), and $\gamma$ is a lower bound on the speed at which interference decays with distance. By Theorem 2, choosing
@@
optimizes the rate of convergence of $\hat{\theta}$. This formalizes how domain knowledge of interference $\gamma$ informs the design of clusters. The right-hand side is increasing in $\gamma$ since less interference means the same level of bias reduction can be achieved with more clusters. It is decreasing in density $\xi_{n}$ because having more units in a given area effectively corresponds to greater interference (an infectious disease may spread more easily).
Choosing the number of clusters according to (7) balances the asymptotic orders of the bias and variance of $\hat{\theta}$. However, as discussed in § 4, validity of the CI
$\hat{\theta}\pm 1.96\cdot\hat{\sigma}k^{-1/2}$ (8)
requires the bias to be of smaller order than the variance. This requires an undersmoothed design where $k$ is of smaller order than (7).
We next recommend an undersmoothed choice of $k$ given domain knowledge of $\gamma$. Let $\mathcal{V}$ denote the area or volume of the study region containing $\mathcal{N}_{n}$. As discussed below, $\mathcal{V}$ is of order $n/\xi_{n}$, so the rate-optimal formula (7) can be rewritten as $k\sim\mathcal{V}^{\frac{2\gamma}{2\gamma+d}}$. While one could choose $k$ equal to the right-hand side, this choice would be extremely sensitive to the unit of length used to measure geographic distance. Switching from square kilometers to square millimeters would dramatically increase $\mathcal{V}$.
Our observation is that both the unit of length and $\gamma$ determine the speed at which interference decays with distance. For any given choice of $\gamma$, the rate of decay is significantly faster if distance is measured in millimeters compared to kilometers, for example. Therefore, domain knowledge of interference informs both $\gamma$ and the unit of length, and one can determine $k$ from these ingredients as follows.
@@
The next subsection provides empirical examples of calibrating $k$ under the conservative choice $\tilde{\gamma}=d$. The last subsection discusses how to determine $\gamma$ to obtain less conservative estimates.
@@ 5.1 Empirical Examples
Sur et al. (2009) conduct a CRT in an urban slum in India spanning about 1.2 by 0.7 km with 38k participants. To compute our suggested number of clusters (9), we need to select $\tilde{\gamma}$ and the unit of length. Suppose we choose the conservative bound $\tilde{\gamma}=d=2$, so that interference is assumed to decay like $r^{-2}$ or faster with each unit of length $r$. If we take 35m to be the unit of length, then for units $i,j,k$ such that $\rho(i,j)=35$m and $\rho(i,k)=70$m, the extent to which $k$’s treatment affects $i$ is less than $1/4$th ($0.252^{-\tilde{\gamma}}=$ 0.25) as small as the extent to which $j$’s affects $i$. This is with only a 35m difference in distance. For this unit of length, (9) yields $k=78$. In other words, these are the assumptions on interference that justify the authors’ choice of $k=80$, originally determined by a conventional power analysis based on a partial interference model.
@@
Next consider Homan et al. (2016) whose trial region is substantially larger at roughly 12 by 4 km with nearly the same number of units (34k). Due to the lower density, there is less bias from interference, so a choice of $k=81$ can be justified under weaker assumptions on interference. Specifically, if $\tilde{\gamma}=2$ but the unit of length is now in 250m increments, then (9) results in $k=84$. For context, the Anopheles mosquito that is the subject of their trial typically does not fly more than 2 km from their breeding grounds (CDC, 2025).
@@ 5.2 Bounding Interference
Optimal design generally requires prior information on certain population parameters. The standard power analysis for CRTs assumes partial interference and requires knowledge of the intracluster correlation coefficient, a measure of within-cluster dependence in potential outcomes (e.g. Baird et al., 2018; Hemming et al., 2017). Our results suggest that, if cross-cluster interference is of first-order importance, the focus of attention should instead be $\gamma$. As illustrated in the previous subsection, the conservative choice $\gamma=d$ can result in relatively small values of $k$ because it allows for a greater degree of interference, so to the extent that one can justify stronger assumptions on interference, that is, values of $\gamma$ that are larger than $d$ for a given unit of length, this would improve asymptotic power.
In the context of infectious disease trials, Halloran et al. (2017) argue that CRT design should be informed by simulating models of disease transmission. Several papers utilize parametric models and simulation methods to estimate or bound contamination bias. Alexander et al. (2020) and Jarvis et al. (2019) use spatial models to provide evidence of contamination in prior CRTs. Multerer et al. (2021) employ models of disease transmission for a similar purpose. Our theory provides a precise way in which modeling can inform design, namely by providing plausible values of $\gamma$. This relates to Staples et al. (2015) who show how to estimate a different measure of cross-cluster interference to assess the degree to which the conventional power analysis overstates trial power.
@@ 6 Simulation Study
Let $\{\tilde{\varepsilon}_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(-0.5,1)$, $\{\beta_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(2,1)$, and $\{\gamma_{i}\}_{i\in\mathcal{N}_{n}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(1,1)$ be independent and drawn independently of locations. We generate spatially autocorrelated errors $\varepsilon_{i}=\tilde{\varepsilon}_{i}+\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)\leq 1\}\tilde{\varepsilon}_{j}/\sum_{k\in\mathcal{N}_{n}}\bm{1}\{\rho(i,k)\leq 1\}$. For $w_{ij}=\min\{\rho(i,j)^{-5},1\}$, we generate outcomes according to
$Y_{i}=\sum_{j\in\mathcal{N}_{n}}w_{ij}D_{j}\beta_{j}+\sum_{j\in\mathcal{N}_{n}}w_{ij}D_{i}D_{j}\gamma_{j}+\varepsilon_{i}.$
@@
We present results for the indirect and total effects using 5000 simulation draws where within each draw, we redraw potential outcomes and recompute the design-based estimand. In Table 1, the “Spatial Interference” columns correspond to the outcome model described above, whereas the “Partial Interference” columns redefine $w_{ij}=0$ if $i,j$ lie in different clusters to eliminate cross-cluster interference. The “CI” rows report the coverage of 95-percent CIs using the indicated standard errors. For our estimator, the “SE” row corresponds to standard errors obtained from our variance estimator (4), while for difference in means, it corresponds to conventional cluster-robust SEs. The “SE” rows are the true superpopulation standard errors obtained by taking the standard deviation of the estimator across the simulation draws. As such, SE should be consistent for SE rather than conservative, while the CIs have asymptotically conservative coverage for the design-based estimand. Finally, row “Excl” is the percentage of units that are not well surrounded.
@@
Table 2 presents results for our estimator under spatial interference, but we multiply (2) by $c\in\{0.8,1.2\}$. This is to assess robustness and illustrate a bias-variance trade-off in the choice of $r_{n}$. The $c=0.8$ columns show that this results in about half the proportion of excluded units relative to Table 1. As a result, the bias is higher, resulting in undercoverage. The $c=1.2$ columns show that the proportion of excluded units a little less than doubles. The bias is lower, and as a result, the probability of coverage is higher. On the other hand, the variance increases, as can be seen in the standard error columns.
@@ 7 Empirical Application
We apply our estimator to data from the unconditional cash transfer experiment mentioned in Example 2. In the experiment, households eligible for the transfer (the treatment) live in homes with thatched roofs, which is a proxy for poverty. Houses are grouped in villages, which are grouped in “sublocations.” See Figure A.2 of Egger et al. (2022) for a map of the study area. The CRT randomizes sublocations (the clusters) into treatment with probability $q=0.5$. Within treatment (control) sublocations, villages are randomized into treatment with probability $p_{1}=2/3$ ($p_{0}=1/3$). Within treated villages, all eligible households receive cash transfers totaling 1000 USD, which is about 75 percent of average annual household spending. Sublocations contain 7.8 villages on average (SD 3.9).
@@
However, the authors note the potential for interference across sublocations (their quote in Example 2). For this reason, their preferred specification is the following “IV” (instrumental variables) regression. The main regressors are the amount of cash transferred per capita to the household’s village and the amount transferred to neighboring villages within a band of $r-2$ to $r$ km from the ego’s village for a range of $r$ values. The corresponding instruments are respectively an indicator for the ego’s village being treated and the share of eligible households in the band assigned to treatment. They use a BIC criterion to select the maximum range of $r$, which is 2 km. Using these estimates, the authors compute a model-assisted estimate of the total effect $\theta^{*}(1,0,2/3,0)$. This compares saturation levels of two-thirds and zero, which is nonparametrically unidentified since control villages have a saturation level of one-third. We instead report our estimates of the total effect $\theta^{*}(1,0,2/3,1/3)$.
Table 3 reports the results for a subset of the outcomes, and Table SA.4.1 in the supplementary appendix reports the remainder. Both tables choose $r_{n}$ in $\hat{\theta}$ according to (2), which results in $r_{n}=1.6$ and 39.66 percent of units not well surrounded. In § SA.4 we describe how we construct cluster radii $R_{j}$ used in this formula. The $\hat{\theta}^{+}$ columns correspond to difference-in-means estimates with clustered standard errors.
We find that the difference-in-means estimates of the direct and total effects are comparable to the RF and IV estimates, respectively, despite the distinctions outlined above. Our estimators find larger direct and total effects. Decomposing the total effect into direct and indirect effects, we find that the latter are substantially smaller in magnitude with large standard errors. Compared to difference in means, our estimates tend to be larger in magnitude but with larger standard errors due to the restriction to well-surrounded units.
@@
To estimate spillover effects, Egger et al. (2022) rerun their IV specification using only non-eligible households, which did not receive any transfers (their §3.3). In Table SA.4.3 of the supplementary appendix, we compare their results with design-based estimates of the overall effect $\theta^{*}(\varnothing,\varnothing,2/3,1/3)$ on the population of non-eligibles. The effect sizes of our estimators and those of difference in means are fairly similar in magnitude to their IV results, but the standard errors are large. Combined with the results in Table 3, we ultimately find strong direct effects of the cash transfers but weaker evidence for spillover effects compared to Egger et al. (2022).
@@ 8 Conclusion
When interference occurs across clusters, conventional analyses of CRTs suffer from bias induced by units near cluster boundaries. To reduce bias at the analysis stage, we provide in § 3.2 an estimator $\hat{\theta}$ that improves upon the fried-egg design by excluding from estimation units that are not surrounded by clusters assigned to the same treatment arm. This may be employed as a robustness check for difference in means. To reduce bias at the design stage, we propose a rate-optimal formula for the number of clusters $k$ in (9). Unlike the standard power analysis that assumes partial interference, our choice balances power against the need to reduce bias and accounts for the density of units in the spatial region. Given $k$, we suggest automating cluster construction using $k$-medoids and prove that the resulting clusters are balanced and globular, thereby approximately minimizing the number of units near boundaries. We also provide valid design-based standard errors.
Under the conventional superpopulation, partial interference framework, power calculations for choosing $k$ require prior knowledge of the intracluster correlation coefficient (ICC). Under our design-based framework, the optimal choice of $k$ requires knowledge of $\gamma$, the speed at which interference decays with distance, rather than the ICC. Absent domain knowledge, one can conservatively set $\gamma$ to the dimension of the spatial region. We discuss in § 5.2 how to obtain less conservative estimates via modeling or prior data.
@@
Whereas $k$-means allows centroids to be any element of $\mathcal{X}$, $k$-medoids constrains the centroids to the data $\mathcal{N}_{n}\subseteq\mathcal{X}$, but in practice the algorithms tend to produce similar output. Our theoretical results pertain to the implementation in Algorithm SA.1.1, the “partitioning around medoids” algorithm. This is simple to understand, though it has complexity $O(k(n-k)^{2})$ compared to an $O(n^{2})$ runtime under the fastest known implementation (Schubert and Rousseeuw, 2021).
@@
Call the density-homogeneous subregions $\mathcal{S}_{1},\ldots,\mathcal{S}_{m}$. For each $\mathcal{S}_{j}$ the researcher can compute the optimal number of clusters $k_{j}$ given in (9), subdivide $\mathcal{S}_{j}$ into $k_{j}$ clusters $C_{j1},\ldots,C_{jk_{j}}$, say using $k$-medoids, and cluster-randomize across the collection of all clusters $\{C_{j\ell}\colon\ell=1,\ldots k_{j},j=1,\ldots,m\}$. Lastly, we recommend modifying the formula for $r_{n}$ since cluster radii may vary substantially across subregions with differing densities. In the definition of $T_{ti}$, replace $r_{n}$ with half the median cluster radius among clusters in the subregion containing $i$.
@@
We next discuss how $\hat{\sigma}^{2}$ relates to variance estimators in the literature. Define $\bm{A}(1)$ ($\bm{A}(2)$) as the symmetric, $n\times n$ matrix with $ij$th entry $A_{ij}(1)$ ($A_{ij}(2)$), and recall that $\hat{\sigma}^{2}(2)$ is the cluster-robust variance estimator, while $\hat{\sigma}^{2}(1)$ is analogous to the Leung (2022b) variance estimator. The proof of Theorem 5 shows that the cluster-robust variance estimator can be asymptotically decomposed as $\hat{\sigma}^{2}(2)=\sigma_{n}^{2}+\mathcal{B}_{n}+o_{p}(1)$ where
@@
The proof further shows that $\hat{\sigma}^{2}(1)=\sigma_{n}^{2}+\mathcal{B}_{n}(1)+o_{p}(1)$ where $\mathcal{B}_{n}(1)$ is obtained by replacing $\bm{A}(2)$ with $\bm{A}(1)$ in (SA.3.1). This replacement adds nonzero off-diagonal elements to $\bm{A}(2)$, so $\bm{A}(1)$ is not block-diagonal and hence not guaranteed to be positive semidefinite. Accordingly, Leung imposes additional conditions to show that $\mathcal{B}_{n}(1)\stackrel{{\scriptstyle p}}{{\longrightarrow}}c\geq 0$ in the superpopulation. We find, however, that $\mathcal{B}_{n}(1)=\mathcal{B}_{n}+o_{p}(1)$ without any conditions on the superpopulation, so in fact $\hat{\sigma}^{2}(1)$ is asymptotically conservative in a purely design-based setup. Faridani and Niehaus (2024) provide a different approach to conservative inference for the GATE based on bounding $\mathcal{B}_{n}$.
The positive-semidefiniteness of the “kernel” $\bm{A}(2)$ results in both a non-negative variance estimator for any $n$ and asymptotic conservativeness. This mirrors the corresponding insight for HAC variance estimators, that positive semidefinite kernels ensure conservativeness in finite-population models. This fact was first pointed out by Leung (2019) and has since been exploited by other papers. (In his case of network-dependent data, the difficulty with using positive-semidefinite HAC kernels is that they are sloped and tend to severely over-reject in finite samples, which is why Leung (2022a) recommends use of the uniform kernel even though it is not positive semidefinite. For spatial data, a variety of positive semidefinite HAC kernels exist, and these tend to have less severe issues with over-rejection relative to the network case.)
Hudgens and Halloran (2008) propose variance estimators for difference-in-means type estimators under partial interference. Their theory relies on an additional stratified interference assumption, which says that potential outcomes only depend on the ego’s treatment assignment and the proportion of treated units in the ego’s cluster, in which case $Y_{i}=Y_{i}(D_{i},p)$ (since they assume complete randomization). Because $p$ is a constant, the dependence structure is the same as under no interference since outcomes within a cluster are only correlated if treatment assignments are. The variance estimators proposed by Hudgens and Halloran (2008) heavily rely on this structure. In contrast, we do not impose partial, let alone stratified, interference, so our setting features both within- and cross-cluster dependence in outcomes. It is therefore critical to account for additional covariance terms that are absent in the Hudgens and Halloran (2008) variance formula to avoid anti-conservativeness. Our estimator does so through the terms involving $A_{ij}(u)$ for $i\neq j$.
@@
We bound the discrepancy between $\tilde{Z}_{i}$ and $X_{i}^{r}$. For any $r\geq r_{n}$ and $t\in\{0,1\}$, $T_{ti}$ is measurable with respect to the $\sigma$-algebra generated by $\mathcal{F}_{i}(r)$. Then by Assumption 3, for any $q>0$, $r\geq r_{n}$, $n$ sufficiently large, and $t\in\{0,1\}$,
@@
because $T_{ti}^{+}Y_{i}/p_{t}^{+}$ is uniformly asymptotically bounded by Assumption 1 and Lemma SA.5.3, and $p_{t}^{+}-\hat{p}_{t}^{+}\precsim k_{n}^{-1/2}$ by Lemma SA.5.6. It suffices to derive the rate of convergence of $\tilde{\theta}^{+}$.
By the argument used to bound (SA.6.4) in the proof of Theorem 2, $\text{Var}(\tilde{\theta}^{+})\precsim k_{n}^{-1}$. It remains to bound the bias, that is, to show that
@@
$\max_{j}\lvert\mathcal{B}(C_{j})\rvert\leq\sum_{\ell=0}^{\lceil\alpha_{n}\rceil}\lvert\mathcal{N}(m_{j},L_{n}+\ell+1)\backslash\mathcal{N}(m_{j},L_{n}+\ell)\rvert\precsim\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}.$ (SA.6.15)